In the 2D squeezing flow, the plate at y=0 is stationary, and the plate at y=h is given a velocity of u = (0, -V). Investigate the axi-symmetric squeezing flow.
Answer #1

The system is as shown below:

insinite plate y=h TTTTTT yo Lisad plate

Consider a cylindrical control volume (CV) of radius r and length L, as shown below:

e axis of symmetry

The walls of the CV are expanding with a radial velocity of Vw as shown. All the fluid particles at |x|=r will be having same radial velocity.

Since the fluid is incompressible, therefore volume of the CV must be constant.

Vol = volume of CV = 2*pi*r*L _________ eq. 1 (here pi = 22/7)

Therefore, dvold(2 * pi *r* L) = 0    (since volume of CV is constant)

. dr de => L+ + * df dr = 0

or L *VW- r*V = 0 (since dre =Vw and de dt dy )

also L = h - V*t,

=> Vw=r*V/L=r*V/(h - V *t) __________________ eq. 2

Also,

Aw = radial acceleration = dv w dt

dt

(using = Vw dt )

(h - V *t)2 *(h - V *t)2

2*V?*r = (h - V * t)2 ____________________ eq. 3

Therefore, the velocity and acceleration of the fluid particles vary with radius r from axis of symmetry and with time as well.